OpenStudy (smurfy14):

(2x-1)^2|x+3|≤0 Solve the inequality (2x-1)^2|x+3|≤0 Solve the inequality @Mathematics

OpenStudy (mertsj):

It is not possible that this product would be negative since (2x-1)^2 will always be positive or zero and |x+3| will always be positive or zero. But it could be zero is 2x-1^2 = 0 or x+3 = 0. 2x-1 = 0 if x = 1/2 and x+3 = 0 if x = -3. So the two possible answers are 1/2 and -3

OpenStudy (smurfy14):

so the answer would be \[[1/2]\cup[-3]\] ?

OpenStudy (mertsj):

yes although that is a weird way to state it. If you want to use set notation it would be {1/2,-3}

OpenStudy (smurfy14):

ok thanks :)